2. The following table shows data on the number of parties attended by different college majors. Test whether there is a difference in the number of parties attended by conducting a Chi-Square Goodness-of-Fit test. Set alpha α = .05 Engineering English Math Psychology Sociology 2 7 4 7 5 Step 1: State the null and research hypotheses in symbols Step 2: Find the critical value (1 point) Step 3: Compute the appropriate test statistic Step 4: Make a decision about the null Step 5: What is your conclusion 2A. Is there a significant difference in the number of parties attended by different majors? (highlight one) Yes No
Answer #1

n=9 á - $21361 o = $865 95% CI = 0.95 Q81-0.95 -0.os n9 df = n-1 = 9-1=8 tat da, df = to. 025,8 = 12.306 (from t table) 95%CI

5 conclusion : There is insufficient evidence t support the claim that the number of parties attended by different college majors

2A) NO